OpenStudy (anonymous):

don't know if doing right? Use the definition to find the Taylor Series (centered at c ) for the function f(x)=sin(x), c=pie/4

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

ick that dont sound good

OpenStudy (anonymous):

you are writing in terms of \[(x-\frac{\pi}{4})\] right?

OpenStudy (anonymous):

first term is \[\cos(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

actually this one is not so bad since you take the derivative of cosine, get -sine, then -cosine, then sine etc etc

OpenStudy (anonymous):

so next one the derivative is -sine and \[-\sin(\frac{\pi}{4})=-\frac{\sqrt{2}}{2}\] so next term is \[-\frac{\sqrt{2}}{2}x\]

OpenStudy (anonymous):

\[\sqrt{2}\]i got cos(\[\pi/4\]=1/ which equals .707

OpenStudy (anonymous):

take the derivative of - sine, get - cosine and keep going. it is going to be - - + + - - + + etc and you will always get \[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

just use \[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

and don't forget to divide by n!

OpenStudy (anonymous):

1/\[\sqrt{2}\]

OpenStudy (anonymous):

so you get \[\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}x-\frac{\sqrt{2}}{2}\frac{x^2}{2!}+\frac{\sqrt{2}}{2}\frac{x^3}{3!}+\frac{\sqrt{2}}{2}\frac{x^4}{4!}-...\]

OpenStudy (anonymous):

you can use \[\frac{1}{\sqrt{2}}\] if you like

OpenStudy (anonymous):

same number.

OpenStudy (anonymous):

oh man i wrote it wrong i am sorry let me do it correctly!

OpenStudy (anonymous):

\[what i got was sinx=x-(x-c)^3! +(x-c)^5/5!-(x-c)^7/7!\]

OpenStudy (anonymous):

x-(x-c)^3/3!

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}(x-\frac{\pi}{4})-\frac{\sqrt{2}}{2}\frac{(x-\frac{\pi}{4})^2}{2!}+\frac{\sqrt{2}}{2}\frac{(x-\frac{\pi}{4})^3}{3!}+\frac{\sqrt{2}}{2}\frac{(x-\frac{\pi}{4})^4}{4!}-...\]

OpenStudy (anonymous):

no this one has both even and odd terms \[\cos(x)\] is even but \[\cos(x-\frac{\pi}{4})\] is not

OpenStudy (anonymous):

oh crap you are doing sine and i am doing cosine. i should just quit. doesn't really matter much though

OpenStudy (anonymous):

goes sine cosine -sine -cosine sine cosine etc

OpenStudy (anonymous):

at \[\frac{\pi}{4}\] you get \[\frac{1}{\sqrt{2}}\] \[\frac{1}{\sqrt{2}}\] \[-\frac{1}{\sqrt{2}}\] \[-\frac{1}{\sqrt{2}}\] \[\frac{1}{\sqrt{2}}\] \[\frac{1}{\sqrt{2}}\] etc

OpenStudy (anonymous):

so your answer is \[\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}(x-\frac{\pi}{4})-\frac{1}{\sqrt{2}}\frac{(x-\frac{\pi}{4})^2}{2}\] \[-\frac{1}{\sqrt{2}}\frac{(x-\frac{\pi}{4})^3}{3!}+\frac{1}{\sqrt{2}}\frac{(x-\frac{\pi}{4})^4}{4!}\]etc

OpenStudy (anonymous):

two pluses followed by two minus and always over n!

OpenStudy (anonymous):

ok so I see if I can replace x with -x to see if even or not. I think this is what is confusing me with most problems. I don't know how to tell what will or should end up in my answer

OpenStudy (anonymous):

not sure how to tell even odd or both on most questions